# The Diophantine problem in Thompson’s group $F$

### v2

@@ 3. Proof of main theorem
and thus $\operatorname{Mon}\langle x_{0}x_{1}^{k}\rangle$ is equationally definable. The case of $k=0$ gives that $\operatorname{Mon}\langle x_{0}\rangle$ is equationally definable, and thus we have shown (3).
@@
    $\{(g,h)\in F\times F\mid\operatorname{expsum}_{x_{0}}(g)=\operatorname{expsum}_{x_{1}}(h)\}$.
@@
Finally, to show (5), the set $\{(g,h)\in F\times F\mid\operatorname{expsum}_{x_{0}}(g)=\operatorname{expsum}_{x_{1}}(h)\}$ is equal to
@@
3.2, (3) and (4) thus now show that this set are all equationally definable. ∎
@@ Acknowledgements
The authors would like to thank James Belk and Corentin Bodart for pointing out that $F$ is isomorphic to a centraliser in $T$, and Collin Bleak, Corentin Bodart and Matthew Brin for finding and explaining references. Elliott was supported by the Heilbronn Institute for Mathematical Research. Levine was supported by the EPSRC Fellowship grant EP/V032003/1 ‘Algorithmic, topological and geometric aspects of infinite groups, monoids and inverse semigroups’.

### v3

@@ 3. Proof of main theorem
and thus $\operatorname{Mon}\langle x_{0}x_{1}^{k}\rangle$ is equationally definable. In particular, we have shown (5). The case of $k=0$ gives that $\operatorname{Mon}\langle x_{0}\rangle$ is equationally definable, and thus we have shown (3).
@@
    $\{(g,h)\in F\times F\mid\operatorname{expsum}_{x_{0}}(g)=\operatorname{expsum}_{x_{1}}(h)\geq 0\}$.
@@
Finally, to show (5), the set $\{(g,h)\in F\times F\mid\operatorname{expsum}_{x_{0}}(g)=\operatorname{expsum}_{x_{1}}(h)\geq 0\}$ is equal to
@@
3.2 (3) and (4) thus now show that this set is equationally definable. ∎
@@ 4. Further questions
One can restrict further still, and look at quadratic equations: equations where each variable occurs exactly twice. This area has seen a lot of interest recently with a wide variety of positive decidability results, including the Grigorchuk group [LysenokMiasnikovUshakov], solvable Baumslag-Solitar groups [LysenokUshakov], and Lamplighter groups [KharlampovichLopezMyasnikov, UshakovWeiers], amongst many others.
Question 4.4.
Is it decidable whether a quadratic equation in $F$ admits a solution?
@@ Acknowledgements
The authors would like to thank James Belk and Corentin Bodart for pointing out that $F$ is isomorphic to a centraliser in $T$, and Collin Bleak, Corentin Bodart and Matthew Brin for finding and explaining references. The authors would like to thank the anonymous referee for their helpful comments. Elliott was supported by the Heilbronn Institute for Mathematical Research. Levine was supported by the EPSRC Fellowship grant EP/V032003/1 ‘Algorithmic, topological and geometric aspects of infinite groups, monoids and inverse semigroups’.